OpenStudy (anonymous):

Check my answer? I needed to make these equations into y = mx + b form: 42x - 4y = 3, 21x = 2y - 3, and 21x = y + 3/2. My answer for #1: y = - 10.5x - (3/4). For #2: y = 10.5x + 1.5, For #3: y = - 21x + 3/2.

OpenStudy (phi):

you are getting the wrong signs on answers #1 and #3

OpenStudy (anonymous):

What do you mean by signs?

OpenStudy (phi):

42x - 4y = 3 can you do this one step by step ?

OpenStudy (anonymous):

I subtract 42x from both sides and get - 4y = - 42x + 3, then I divide everything by - 4, to isolate y and get y = - 10.5x - (3/4).

OpenStudy (phi):

sounds good, except -42x/-4 is not -10.5x

OpenStudy (phi):

minus over minus makes a positive

OpenStudy (anonymous):

It's positive 10.5, just realized my mistake.

OpenStudy (phi):

that's what I meant by getting the wrong sign

OpenStudy (anonymous):

Okay, thanks.

OpenStudy (phi):

for the last question 21x = y + 3/2 subtact 3/2 from both sides 21x - 3/2 =y or y = 21x -3/2

OpenStudy (anonymous):

Thanks